Solve the system of equations shown below using elimination, and then state the solution as an ordered pair.
4x+5y=-18
-9x+2y=-92

Found 2 solutions by richwmiller, stanbon:
Answer by richwmiller(17219)

multiply the first by 2 and the second by -5 and add them
solve for x
then plug that into any of the equations and solve for y
then write them as (x,y)

Answer by stanbon(75887)

4x+5y=-18
-9x+2y=-92
-------
Multiply thru 1st equation by 2
Multiply thru 2nd equation by 5
-------------------------------------
8x + 10y = -36
-45x + 10y = -460
------------------------
53x = 424
x = 8
----
Solve for "y":
4x + 5y = -18
32 + 5y = -18
5y = -50
y = -10
----
Ans: (8,-10)
==================
Cheers,
Stan H.
==================